# Poisson Midpoint Method for Log Concave Sampling: Beyond the Strong Error Lower Bounds

### v1

@@ 1 Introduction
Sampling from a density $\pi(x)\propto\exp(-F(x))$ over $\mathbb{R}^{d}$ is of fundamental interest in physics, economics, and finance johannes2010mcmc ; von2011bayesian ; kobyzev2020normalizing . Applications in computer science include volume computation vempala2010recent and bandit optimization russo2018tutorial . A popular approach is Langevin Monte Carlo (LMC) which is the Euler-Maruyama discretization of the continuous time Ito Stochastic Differential Equation (SDE) called (overdamped/underdamped) Langevin Dynamics. The convergence of LMC has been extensively studied in the literature durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021convergencelangevinmontecarlo ; cheng2017convergencelangevinmcmckldivergence ; cheng2018-underdamped-lmc ; dalalyan2020sampling under various assumptions on the potential function $F$ such as log concavity and isoperimetry.
The randomized midpoint discretization for Langevin dynamics (RLMC), introduced by shen2019randomizedmidpointmethodlogconcave and developed further in yu2023langevinmontecarlostrongly ; he2021ergodicitybiasasymptoticnormality ; altschuler2024shiftedcompositioniiilocal considers a more sophisticated alternative to Euler-Maruyama discretization, reducing the bias while introducing variance, leading to faster convergence bounds than for LMC. The work of Cao et. al Cao_2021 established lower bounds for the computational complexity for convergence in the strong $L^{2}$ error for a class of sampling algorithms, which includes RLMC and LMC. This demonstrated that the computational complexity of RLMC with underdamped Langevin dynamics is optimal in dimension and accuracy (up to log factors) in terms of the strong $L^{2}$ error.
@@ 2.3 Prior work
Recent works have focused on the rigorous theoretical analysis of classical and popular MCMC algorithms to establish complexity bounds and theoretical limits. The prototypical case of Overdamped LMC has been studied when the target $\pi$ is strongly log-concave and more generally $\pi$ when satisfies isoperimetric inequalities dalalyan2017theoretical ; durmus2017unadjusted ; durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021-tail-growth ; mou2022improved ; sinho-non-log-concave-lmc . Underdamped LMC has been considered as a faster alternative. This case too has been well studied when $\pi$ is strongly log-concave and when $\pi$ satisfies isoperimetric inequalities cheng2018-underdamped-lmc ; dalalyan2020sampling ; ganesh2020faster ; ma2021there ; zhang2023improveddiscretizationanalysisunderdamped
LMC is the Euler-Maruyama discretization of continuous time Langevin Dynamics, which can lead to sub-optimal convergence due to statistical bias in the approximation. Thus, shen2019randomizedmidpointmethodlogconcave introduced the randomized midpoint method for LMC (RLMC) which reduces the bias in the approximation by introducing a randomized estimator at the cost of higher variance. RLMC does not involve higher order derivatives of $\nabla F$ as in Runge-Kutta schemes for SDEs kloeden1992stochastic - allowing its use for generative modeling with denoising diffusion models kandasamy2024poissonmidpointmethodlangevin . This leads to improvement in the convergence rates compared to LMC under log concavity (see Table 1). he2021ergodicitybiasasymptoticnormality ; yu2023langevinmontecarlostrongly sharpen the bounds in shen2019randomizedmidpointmethodlogconcave .
@@ 3 Results
Our assumption on the initialization is very standard vempala2022rapidconvergenceunadjustedlangevin ; shen2019randomizedmidpointmethodlogconcave , and satisfied (for example) by $\mathcal{N}(x^{*},\mathbf{I}_{d}/L).$ With $p=3$, we get a complexity of $\tilde{O}(\frac{\kappa^{7/6}d^{1/6}+\kappa^{13/10}d^{1/5}}{\epsilon^{1/3}}).$ The complexity bound for Underdamped LMC is cheng2018-underdamped-lmc $\tilde{O}(\kappa^{3/2}/\epsilon)$, and that of Underdamped RLMC is $\tilde{O}(\frac{\kappa^{7/6}}{\epsilon^{1/3}}+\frac{\kappa}{\epsilon^{2/3}})$. To our knowledge, PLMC is the best known discretization of underdamped Langevin dynamics in terms of $\epsilon$.

### v2

@@ 1 Introduction
Sampling from a density $\pi(x)\propto\exp(-F(x))$ over $\mathbb{R}^{d}$ is of fundamental interest in physics, economics, and finance johannes2010mcmc ; von2011bayesian ; kobyzev2020normalizing . Applications in computer science include volume computation vempala2010recent and bandit optimization russo2018tutorial . A popular approach is Langevin Monte Carlo (LMC) which is the Euler-Maruyama discretization of the continuous time Ito Stochastic Differential Equation (SDE) called (overdamped/underdamped) Langevin Dynamics. The convergence of LMC has been extensively studied in the literature durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021convergencelangevinmontecarlo ; cheng2017convergencelangevinmcmckldivergence ; cheng2018-underdamped-lmc ; dalalyan2020sampling ; altschuler2025shiftedcompositionivunderdamped under various assumptions on the potential function $F$ such as log concavity and isoperimetry.
The randomized midpoint discretization for Langevin dynamics (RLMC), introduced by shen2019randomizedmidpointmethodlogconcave and developed further in yu2023langevinmontecarlostrongly ; he2021ergodicitybiasasymptoticnormality ; altschuler2024shiftedcompositioniiilocal ; altschuler2025shiftedcompositionivunderdamped considers a more sophisticated alternative to Euler-Maruyama discretization, reducing the bias while introducing variance, leading to faster convergence bounds than for LMC. The work of Cao et. al Cao_2021 established lower bounds for the computational complexity for convergence in the strong $L^{2}$ error for a class of sampling algorithms, which includes RLMC and LMC. This demonstrated that the computational complexity of RLMC with underdamped Langevin dynamics is optimal in dimension and accuracy (up to log factors) in terms of the strong $L^{2}$ error.
@@ 2.3 Prior work
Recent works have focused on the rigorous theoretical analysis of classical and popular MCMC algorithms to establish complexity bounds and theoretical limits. The prototypical case of Overdamped LMC has been studied when the target $\pi$ is strongly log-concave and more generally $\pi$ when satisfies isoperimetric inequalities dalalyan2017theoretical ; durmus2017unadjusted ; durmus2018analysislangevinmontecarlo ; vempala2022rapidconvergenceunadjustedlangevin ; erdogdu2021-tail-growth ; mou2022improved ; sinho-non-log-concave-lmc . Underdamped LMC has been considered as a faster alternative. This case too has been well studied when $\pi$ is strongly log-concave and when $\pi$ satisfies isoperimetric inequalities cheng2018-underdamped-lmc ; dalalyan2020sampling ; ganesh2020faster ; ma2021there ; zhang2023improveddiscretizationanalysisunderdamped ; altschuler2025shiftedcompositionivunderdamped
LMC is the Euler-Maruyama discretization of continuous time Langevin Dynamics, which can lead to sub-optimal convergence due to statistical bias in the approximation. Thus, shen2019randomizedmidpointmethodlogconcave introduced the randomized midpoint method for LMC (RLMC) which reduces the bias in the approximation by introducing a randomized estimator at the cost of higher variance. RLMC does not involve higher order derivatives of $\nabla F$ as in Runge-Kutta schemes for SDEs kloeden1992stochastic - allowing its use for generative modeling with denoising diffusion models kandasamy2024poissonmidpointmethodlangevin . This leads to improvement in the convergence rates compared to LMC under log concavity (see Table 1). he2021ergodicitybiasasymptoticnormality ; yu2023langevinmontecarlostrongly ; altschuler2024shiftedcompositioniiilocal ; altschuler2025shiftedcompositionivunderdamped sharpen the bounds in shen2019randomizedmidpointmethodlogconcave .
@@ 3 Results
Our assumption on the initialization is very standard vempala2022rapidconvergenceunadjustedlangevin ; shen2019randomizedmidpointmethodlogconcave , and satisfied (for example) by $\mathcal{N}(x^{*},\mathbf{I}_{d}/L).$ With $p=3$, we get a complexity of $\tilde{O}(\frac{\kappa^{7/6}d^{1/6}+\kappa^{13/10}d^{1/5}}{\epsilon^{1/3}}).$ The complexity bound for Underdamped LMC is cheng2018-underdamped-lmc $\tilde{O}(\kappa^{3/2}/\epsilon)$, and that of Underdamped RLMC is $\tilde{O}(\frac{\kappa^{7/6}}{\epsilon^{1/3}}+\frac{\kappa}{\epsilon^{2/3}})$ shen2019randomizedmidpointmethodlogconcave . The concurrent work of altschuler2025shiftedcompositionivunderdamped claims a rate of $\tilde{O}(\kappa^{5/6}d^{5/3}/\epsilon^{2/3})$ to achieve $\text{KL}\leq\epsilon^{2}.$ This is in the low friction regime $\gamma\asymp\sqrt{\alpha},$ and for a double midpoint implementation of Underdamped RLMC. To our knowledge, PLMC is the best known discretization of underdamped Langevin dynamics in terms of $\epsilon$.